OpenStudy (acarpenter2):

In the formula D=R*T what relationship exists between the rate and time? Medal and Fan for the first to give me a correct answer

OpenStudy (welshfella):

D = R*T divide both sides by T R = D/T as R increases T decreases so there is an inverse relationship between Rate and Time.

OpenStudy (welshfella):

we say R varies inversely with T

OpenStudy (acarpenter2):

Thanks!

OpenStudy (welshfella):

yw